OpenStudy (anonymous):

Write the equation y - 2 = 4(x + 5) in standard form.

OpenStudy (anonymous):

standard form of a linear equation would be in the format Ax +By = C So over here you just solve the parentheses and rearrange the equation so it's in the above mentioned format Hope this helps!

OpenStudy (anonymous):

i think its -4x+y=22 im not sure

OpenStudy (anonymous):

i did this mentally

OpenStudy (anonymous):

these are the suggestions A. 4x + y = 7 B. 4x - y = 3 C. 4x - y = -22

OpenStudy (anonymous):

c